A gas occupies a volume of 300cm cube at 410mmHg. What will be its pressure when volume is decreased to 130cm cube at constant temperature.
1
Expert's answer
2020-04-18T06:55:47-0400

According to the conditions of the question we should make up two equations of Mendeleev-Clapeyron, which characterize the initial and final state for this gas:

"p(1)•V(1) = \nu•R•T"

"p(2)•V(2)=\nu•R•T"

Since the left sides of the equations are equal to one product ("\nu•R•T"), therefore they are equal to each other:

"p(1)\u2022V(1)=p(2)\u2022V(2)" "\to" "p(2) = (p(1)\u2022V(1))\/V(2)" = (300(cm3)•410(mmHg))/130(cm3) = 946 mmHg

Answer: p(2)=946 mmHg